OpenStudy (anonymous):

The next one is a little tricky. Suppose f(x) = |x|x Then f^{-1}(x)= what? if x\geq0, and f^{-1}(x)=what? if x\leq 0.

hartnn (hartnn):

definition of |x| \(\large |x| =x \quad x\ge0 \\ \large |x|=-x \quad x\le 0 \)

hartnn (hartnn):

so what is f(x) when x>= 0 ?

OpenStudy (anonymous):

i have no idea

hartnn (hartnn):

f(x) = |x|x when x > =0, |x| = x so when x>=0 f(x) = x*x = x^2 makes sense ?

hartnn (hartnn):

yes/no/idk ?

OpenStudy (anonymous):

kind of. so when x is less then 0 it is -x^2?

hartnn (hartnn):

absolutely correct!

OpenStudy (anonymous):

okay i think i get it now. thanks again! may have another one for ya haha

hartnn (hartnn):

welcome ^_^ sure! :)

OpenStudy (anonymous):

question on this one still. I can't figure out what the inverse is for the less then 0 one. if f(x)=-x^2 then wouldnt the inverse function be f-1(x)=sqrt(-x)?

hartnn (hartnn):

y = -x^2 x = -y^2 -x = y^2 \(f^{-1}(x) = \sqrt {-x}\) yes

OpenStudy (anonymous):

thats what i thought but when i plug it in it is wrong.

hartnn (hartnn):

let me think..

hartnn (hartnn):

since -x = |x| when x<=0 try \(f^{-1}(x)=\sqrt{|x|}\)

OpenStudy (anonymous):

still wrong :(

hartnn (hartnn):

:O i would take a last attempt to this since x is negative, we take the negative square root \(f^{-1}(x)= -\sqrt{-x}\) if thats also incorrect, then even idk whats correct

OpenStudy (anonymous):

that was one i tried before, still wrong. :/ thank you though i got the first part of it right!

hartnn (hartnn):

hmm. sorry and welcome ^_^